[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Aug 15, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Aug 15 20:28:19 UTC 2017


HSU Shift Log for Tue Aug 15 2017 16:28:13 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: The ATCF would not let us start genesis area 21 in the  
Atlantic. Contacted Monica, who found files for 21 already in the system.  
She was able to delete them, but there also appeared to be an issue with  
the 70 series not being used sequentially (A21 should have used AL77, but  
that was already taken by Gert for some reason). After Monica fixed it, we  
were able to start A21 using AL78.

N-AWIPS issues: None

AWIPS-2/NCP issues: None

Other Techincal issues: The SDM reported that they have had problems  
hearing us on the phone for the past several days, with a lot of  
interference that has made it difficult for them to understand what we are  
saying. Opened an NCO trouble ticket with priority 2. Followed up later  
with SDM and the phone lines were clear. Marc L plans to replace Pac/Atl  
ops phones.

Miscellaneous issues: GOES-16 meso request for Gert from 21-06 Z
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170815/8908c6d8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list